OpenStudy (emilyjones284):

Using u-sub -5x/(1-2x^2)^2 dx

OpenStudy (danjs):

let u = 1- 2x^2

OpenStudy (danjs):

du = -4x dx

OpenStudy (danjs):

so dx = du / (-4x)

OpenStudy (danjs):

\[-5 \int\limits \frac{ x }{ u }\frac{ du }{ -4x }\]

OpenStudy (emilyjones284):

And I plugged them in in terms of u but then I got confused

OpenStudy (emilyjones284):

Ohhh I didn't factor out the 5

OpenStudy (emilyjones284):

Why isn't it u^2

OpenStudy (danjs):

oh yeah

OpenStudy (danjs):

\[\frac{ 5 }{ 4 } \int\limits u ^{-2}du\]

OpenStudy (emilyjones284):

So it should be u^2?

OpenStudy (danjs):

yeah it is 1 over u^2 or just u^-2 like i wrote... i erased the wrong one and rewrote it

OpenStudy (danjs):

\[\frac{ 5 }{ 4 }*\frac{ -1 }{ u }\]

OpenStudy (danjs):

sub back in u = 1-2x^2